# Exploring the Roles of Large Language Models in Reshaping Transportation Systems: A Survey, Framework, and Roadmap

### v1

@@ 1.1 Motivation
Given these desirable properties of LLMs, the potential of LLMs to address the above challenges of modern ITS has recently gained great popularity. Far beyond conversational chatbots, LLMs can be a transformative force for transportation systems (Qu et al., 2023; Lv, 2023; Mahmud et al., 2025). For example, LLMs improve the accuracy of traffic prediction by integrating spatiotemporal patterns with semantic context (Ren et al., 2024), automate multimodal traffic scenario understanding (Zhou et al., 2023), optimize real-time traffic control through natural language interfaces (Lai et al., 2023), generate high-fidelity simulation environments (Zhao et al., 2024), facilitate communication between vehicles and users (Cui et al., 2024), and hence significantly improve the operational efficiency of ITS. These pioneering studies have suggested that LLMs can play a powerful role in shaping future transportation systems in various aspects and provide promising solutions to address the limitations of current ITS.
@@
Survey Year Venue Taxonomy Scope Focus
Zhou et al. (2023) 2023 IEEE TIV Tasks Perception, scene understanding, navigation, and decision making of autonomous driving Applications of VLMs in main autonomous driving tasks
Yan and Li (2023) 2023 ArXiv Tasks Traffic perception, traffic prediction, traffic simulation, and traffic decision-making Applications of generative AI on different tasks in ITS
Zhou et al. (2024a) 2024 TRANS. RES. E Applications and architecture components Transport plan, network design, management, infrastructure construction in urban mobility systems A conceptual framework that introduces a hierarchically unified generative foundation model
Zhang et al. (2024a) 2024 IJCRS Applications and tasks Traffic management, transportation safety, and autonomous driving The potential of LLMs to increase forecasting accuracy, improve safety, and enhance decision making
Wandelt et al. (2024) 2024 Appl. Sci. Applications Autonomous driving, safety, tourism, traffic, and others Overview of current literature, challenges and research recommendations
Gan et al. (2024) 2024 Adv. Eng. Inform. Tasks and deployment techniques Perception, prediction, control, and simulation for ITS; navigation, planning, and decision making for AV Application and deployment of large models for ITS and AV
Zhang et al. (2024g) 2024 ArXiv Data processing pipelines and models Traffic and demand forecasting Tokenization, prompt, embedding and fine-tune techniques, zero-shot/few-shot prediction
Choi et al. (2024) 2024 ArXiv Models Traffic data generation, traffic estimation and prediction, and unsupervised representation learning Generative models in transportation, such as GAN, VAE, and diffusion models
Mahmud et al. (2025) 2025 IEEE TITS Applications Traffic prediction, signal optimization, traffic control, public transport, and V2X communication, Potential of LLMs in optimizing ITS
This survey Roles of LLMs from a methodological perspective Sensing, learning, modeling, and managing of transportation systems Different roles of LLMs in integration with transportation systems
@@ 1.2 Scope and focus
We are also aware of several related review articles on LLMs or generative AI techniques for transportation in Tab. 1. Our survey differentiates itself from others with a fresh perspective, novel taxonomy, broader coverage, and practical guidance. Furthermore, to support future development and encourage community engagement, we curate a GitHub repository at link: https://github.com/tongnie/awesome-llm4tr, with a supporting collection of datasets, benchmarks, and tools for LLM-driven transportation research. We will continue to update this online project so as to provide a platform for tracking the latested advances in the field. Finally, we hope that this review will provide the research community with a broader perspective that will advance the application of LLMs in transportation research and practices.
@@ 1.3 Contribution
This survey makes four key contributions to the community:
@@ 2.1.2 Learning
Learning refers to pattern recognition and predictive analytics of traffic data. Learning in transportation systems involves both machine learning and data-driven predictive learning approaches that extract actionable insights from large and heterogeneous datasets, bridging the gap between sensing and decision-making (Kumar and Raubal, 2021; Shaygan et al., 2022). Early work focused on statistical methods and shallow machine learning models to assist in pattern recognition and traffic data mining; however, the last decade has witnessed a surge in deep learning, reinforcement learning, and generative learning frameworks (Veres and Moussa, 2019; Haydari and Yılmaz, 2020; Lin et al., 2023). Many complex predictive tasks can be solved by end-to-end learning methods, such as dynamic routing, traffic prediction, anomaly detection, and safety-critical applications (Veres and Moussa, 2019). This shift is driven by the need to handle massive amounts of traffic big data generated by increasing traffic participants and infrastructures with scalability.
Deep learning transforms raw sensor data into actionable insights through three primary paradigms: supervised learning, unsupervised learning, and reinforcement learning (Veres and Moussa, 2019). A representative application in transportation lies in graph-based deep learning, where heterogeneous data (e.g., traffic flow, social media activity, traffic participants) are structured into spatio-temporal graphs to uncover latent correlations and interactions (Xue et al., 2025). Graph neural networks (GNNs) have proven effective for tasks such as traffic prediction and interaction modeling by encoding road network topology and traffic dynamics (Rahmani et al., 2023). In addition, unsupervised learning frameworks (e.g., deep generative models (Choi et al., 2024)) are used to learn the distribution of mobility patterns and generate simulated system states in large-scale networks. More recent advances include physics-informed neural networks that embed traffic flow equations into learning architectures (Shi et al., 2021), multi-agent RL frameworks coordinating connected vehicles or traffic controllers (Farazi et al., 2021), transfer learning techniques enable knowledge sharing across cities with diverse traffic patterns (Tang et al., 2022), and causal learning methods disentangle confounding factors in traffic analysis (Liu et al., 2024f).
@@ 2.1.4 Managing
2.2 Background of Large Language Models
Large language models (LLMs) refer to Transformer-based (Vaswani et al., 2017) language models that contain hundreds of billions (or more) of parameters being trained on massive text data (Zhao et al., 2023; Shanahan, 2024). The Internet-scale training data and extensive model parameters enable LLMs to have impressive capabilities, from a natural language modeler to a general problem solver. To have a quick understanding of modern LLMs, this section briefly introduces the backgrounds and preliminaries of LLMs.
@@ 2.3.3 Post-training
Table 3: Comparison between instruction tuning, supervised fine-tuning, and prompt tuning.
Aspect Supervised Fine-Tuning Instruction Tuning Prompt Tuning
@@
After collecting and constructing feedback datasets from human experts, a prevalent method for alignment is Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019), where models are fine-tuned using datasets of human preferences to guide their behavior. RLHF adapts LLMs to human feedback by learning a reward model, incorporating human in the training loop. It involves three stages:
@@
Parameter-efficient fine-tuning. The discussed instruction tuning and alignment tuning methods typically require full-parameter tuning. Because LLMs are parameter-extensive, it is computationally prohibitive, memory intensive, and risks catastrophic forgetting. Parameter-efficient fine-tuning (PEFT) techniques address these limitations by updating only a small subset of parameters while preserving the model’s inherent capabilities (Liu et al., 2022; Ding et al., 2023). These methods such as adapter tuning, prefix tuning, prompt tuning, and Low-Rank Adaptation (LoRA) strike a balance between task-specific performance and resource efficiency, enabling cost-effective deployment of LLMs across diverse applications. We briefly introduce these them as follows.
@@ 3.1 Conceptual framework
    Excellent reasoning and planning abilities: Chain-of-thought prompting and recursive self-improvement enable LLMs to decompose complex tasks into multi-step reasoning processes with verifiable intermediate states. In addition, LLMs can perform complex tasks by planning to use tools, integrate external knowledge, and learn from demonstrations.
@@
As Figure 1 illustrates, the LLM4TR framework shapes the four-step transportation management cycle through an LLM-centric lens:
@@ 3.2 Taxonomy
Figure 2: The literature classification procedure in this survey.
@@
      Example: Designing reward functions for traffic control agents in reinforcement learning (Yu et al., 2024), assisting in synthesizing virtual driving environments (Zhao et al., 2024), and providing feedback for the refinement of model component (Tian et al., 2024).
@@
Following this structured taxonomy, this survey classifies related literature using some principles shown in Fig. 2. Please note that in some studies the role of LLMs can be multifaceted. For example, they may serve simultaneously as both a context encoder and a decision maker. In our classification, we have identified and assigned each method to its primary role to ensure clarity and consistency.
@@ 3.3 Research trend visualization
As an intuitive overview of the current research trend and focus, we visualize the statistics of selected papers according to our taxonomy in Fig. 3. Several key findings in these figures can be summarized as follows: (1) For the four roles defined in this paper, the "decision facilitator" attracts the most interest in existing studies, where most of them focus on managing and learning tasks such as traffic prediction, autonomous driving, and traffic signal control. This is also one of the most crowded tracks in transportation research. (2) The "component generator" is also a popular direction in which the generative capabilities of LLMs can assist in designing modules or systems. (3) Some areas are still unexplored. For example, adopt LLMs as decision makers in traffic modeling and designers for learning tasks. The former can automate the traffic simulation procedure by acting as a decision engine, while the latter can inspire the neural architecture design for spatial-temporal predictive learning. (4) From another facet, different tasks highlight different functionalities of LLMs. For example, the sensing task populates the integration of multimodal LLMs for information fusion; Among these four tasks, the application of the "knowledge encoder" accounts for the highest proportion in learning. These intuitive visual cues can guide future research choices and systematic overviews of this research field.
@@ 4 LLMs as Information Processors
Due to the in-context learning capability (Brown et al., 2020) and the multimodal extension, LLMs have emerged as versatile tools for processing heterogeneous transportation data, enabling efficient encoding, analysis, and fusion of multimodal information generated by traffic participants. This section explores three key methodological paradigms: (1) contextual encoding of traffic scenarios and task queries, (2) analytical reasoning over structured and unstructured traffic data, and (3) multimodal integration for holistic scenario understanding. These approaches collectively address challenges in handling heterogeneous information, multi-source data, and complex semantics in transportation systems, especially crucial for sensing and learning tasks.
@@ 4.1 Context encoder
A primary application of LLMs lies in encoding textual descriptions of dynamic traffic environments or task queries into machine-readable formats and task-specific patterns. This characteristic can facilitate the problem-solving process of ITS as LLMs can translate natural language instruction to domain-specific context that is compatible with other system components. Studies such as Keysan et al. (2023), Zhao et al. (2024), and Liao et al. (2024) demonstrate how LLMs transform free-form language descriptions of driving scenarios into encodings that can be directly consumed by downstream systems. Advanced implementations extend this capability to traffic simulation frameworks. TP-GPT (Wang et al., 2024a) generates accurate SQL queries for large-scale traffic databases and natural language interpretations by parsing user requests. It also employs a multi-agent collaboration strategy and few-shot learning to handle complex analytical tasks, such as traffic pattern recognition and privacy-preserving data interpretation. TP-GPT outperforms GPT-4 and PaLM 2 on a traffic analysis benchmark called TransQuery. ChatSUMO (Li et al., 2024a) integrates LLMs with the SUMO traffic simulator to democratize traffic simulation for non-experts. The LLM (e.g., Llama3.1) processes natural language inputs and converts them into keywords, which trigger Python scripts to fetch OpenStreetMap data, generate road networks, and configure traffic conditions. By automating code generation and data interpretation, ChatSUMO reduces scenario creation time from 15 minutes to 30 seconds while achieving 96% simulation accuracy.

### v2

@@ 1.1 Motivation
Given these desirable properties of LLMs, the potential of LLMs to address the above challenges of modern ITS has recently gained great popularity. Far beyond conversational chatbots, LLMs can be a transformative force for transportation systems (Qu et al., 2023; Lv, 2023; Mahmud et al., 2025). For example, LLMs improve the accuracy of traffic prediction by integrating spatiotemporal patterns with semantic context (Ren et al., 2024), automate multimodal traffic scenario understanding (Zhou et al., 2023), optimize real-time traffic control through natural language interfaces (Lai et al., 2023), generate high-fidelity simulation environments (Zhao et al., 2024a), facilitate communication between vehicles and users (Cui et al., 2024), and hence significantly improve the operational efficiency of ITS. These pioneering studies have suggested that LLMs can play a powerful role in shaping future transportation systems in various aspects and provide promising solutions to address the limitations of current ITS.
@@
Survey Year Venue Taxonomy Scope Focus Perspective & Contributions
Zhou et al. (2023) 2023 IEEE TIV Tasks Perception, scene understanding, navigation, and decision making of autonomous driving Applications of VLMs in autonomous driving Focuses only on VLMs for AVs without covering broader LLMs’ applications in transportation
Yan and Li (2023) 2023 ArXiv Tasks Traffic perception, prediction, simulation, and decision-making Generative AI applications in ITS tasks Task-based classification; emphasis on general generative models; lacks LLM-specific taxonomy
Zhou et al. (2024a) 2024 TRANS. RES. E Applications and architecture Urban mobility planning and design Hierarchical foundation model concept for urban systems Focuses on urban mobility architecture and models; lacks LLM-specific taxonomy
Zhang et al. (2024a) 2024 IJCRS Applications and tasks Traffic management, safety, autonomous driving LLMs’ benefits in ITS performance metrics Application-level review; without system-level methodological framework
Wandelt et al. (2024) 2024 Appl. Sci. Applications AV, traffic, tourism, etc. Overview of literature and challenges High-level summary of applications; without analytical framework or taxonomy
Gan et al. (2024) 2024 Adv. Eng. Info. Tasks and deployment Perception, prediction, control, and simulation for ITS; AV navigation, planning, and decision making Application and deployment of large models for ITS and AV Task- and deployment-centric; lacks methodological taxonomy of LLMs across system levels
Zhang et al. (2024g) 2024 ArXiv Data pipelines and models Traffic and demand forecasting tasks Tokenization, prompting, embedding and fine-tuning; zero/few-shot prediction Focuses on training/prediction techniques only; lacks system integration perspective
Choi et al. (2024) 2024 ArXiv Models Traffic data generation, estimation, prediction, and representation learning Generative models in transportation, such as GAN, VAE, and diffusion models Emphasis on architectures and applications of deep generative models; not LLM- or role-specific
Mahmud et al. (2025) 2025 IEEE TITS Applications Prediction, control, signal optimization, V2X, and public transport Use of LLMs in ITS optimization Application-centric; lacks methodological or system-level taxonomy; without practical guidance
This survey 2025 AIT Methodological roles of LLMs Broader coverage across transportation domains: sensing, learning, modeling, and managing Functional roles of LLMs in transportation system integration First review to systematically define and categorize LLMs across the transportation lifecycle using a role-centric taxonomy; practical guidelines for role-based integration.
@@ 1.2 Scope and focus
We are also aware of several related review articles on LLMs or generative AI techniques for transportation. Table 1 provides a comprehensive comparison of these surveys in terms of their publication details, taxonomy, scope, focus, and key differences compared to our work. Most of the existing reviews adopt a taxonomy based on application domains or task categories. For example, Yan and Li (2023), Gan et al. (2024) and Mahmud et al. (2025) structure their surveys around application-specific use cases such as traffic prediction, autonomous driving, or traffic management. However, they do not systematically distinguish how LLMs function within transportation systems from a methodological perspective. In contrast, our review differentiates itself from the prevailing tasks- or application-oriented ones with a fresh perspective, novel taxonomy, broader coverage, and practical guidance. We systematically analyze how LLMs can serve distinct methodological functions across the transportation system lifecycle. This taxonomy enables us to unify technical features of LLMs with their practical integration in transportation systems, offering practitioners with actionable guidance on selecting and configuring LLMs for each stage of system design and operation.
Furthermore, to support future development and encourage community engagement, we curate a GitHub repository at link: https://github.com/tongnie/awesome-llm4tr, with a supporting collection of datasets, benchmarks, and tools for LLM-driven transportation research. We will continue to update this online project so as to provide a platform for tracking the latest advances in the field. Finally, we hope that this review will provide the research community with a broader perspective that will advance the application of LLMs in transportation research and practices.
@@ 1.3 Contribution
Our role-centric framework not only unifies disparate application domains but also provides practical considerations that are absent from earlier task- or application-centric reviews. To the best of our knowledge, this is the first review to propose a methodological taxonomy of LLMs in transportation, grounded in their functional roles, and supported by a broad and up-to-date coverage of the literature as well as practical guidelines. Specifically, this survey makes four key contributions to the community:
@@ 2.1.2 Learning
Learning refers to pattern recognition and predictive analytics of traffic data. Learning in transportation systems involves both machine learning and data-driven predictive learning approaches that extract actionable insights from large and heterogeneous datasets, bridging the gap between sensing and decision-making (Kumar and Raubal, 2021; Shaygan et al., 2022). Early work focused on statistical methods and shallow machine learning models to assist in pattern recognition and traffic data mining; however, the last decade has witnessed a surge in deep learning, reinforcement learning, and generative learning frameworks (Veres and Moussa, 2019; Haydari and Yılmaz, 2020; Lin et al., 2023a). Many complex predictive tasks can be solved by end-to-end learning methods, such as dynamic routing, traffic prediction, anomaly detection, and safety-critical applications (Veres and Moussa, 2019). This shift is driven by the need to handle massive amounts of traffic big data generated by increasing traffic participants and infrastructures with scalability.
Deep learning transforms raw sensor data into actionable insights through three primary paradigms: supervised learning, unsupervised learning, and reinforcement learning (Veres and Moussa, 2019). A representative application in transportation lies in graph-based deep learning, where heterogeneous data (e.g., traffic flow, social media activity, traffic participants) are structured into spatio-temporal graphs to uncover latent correlations and interactions (Xue et al., 2025). Graph neural networks (GNNs) have proven effective for tasks such as traffic prediction and interaction modeling by encoding road network topology and traffic dynamics (Rahmani et al., 2023). In addition, unsupervised learning frameworks (e.g., deep generative models (Choi et al., 2024)) are used to learn the distribution of mobility patterns and generate simulated system states in large-scale networks. More recent advances include physics-informed neural networks that embed traffic flow equations into learning architectures (Shi et al., 2021), multi-agent RL frameworks coordinating connected vehicles or traffic controllers (Farazi et al., 2021), transfer learning techniques enable knowledge sharing across cities with diverse traffic patterns (Tang et al., 2022), and causal learning methods disentangle confounding factors in traffic analysis (Liu et al., 2024g).
@@ 2.1.4 Managing
2.2 Background of LLMs
LLMs refer to Transformer-based (Vaswani et al., 2017) language models that contain hundreds of billions (or more) of parameters being trained on massive text data (Zhao et al., 2023; Shanahan, 2024). The Internet-scale training data and extensive model parameters enable LLMs to have impressive capabilities, from a natural language modeler to a general problem solver. To have a quick understanding of modern LLMs, this section briefly introduces the backgrounds and preliminaries of LLMs.
@@ 2.3.3 Post-training
Table 3: Comparison between instruction tuning, SFT, and prompt tuning.
Aspect SFT Instruction Tuning Prompt Tuning
@@
After collecting and constructing feedback datasets from human experts, a prevalent method for alignment is RLHF (Christiano et al., 2017; Ziegler et al., 2019), where models are fine-tuned using datasets of human preferences to guide their behavior. RLHF adapts LLMs to human feedback by learning a reward model, incorporating human in the training loop. It involves three stages:
@@
Parameter-efficient fine-tuning. The discussed instruction tuning and alignment tuning methods typically require full-parameter tuning. Because LLMs are parameter-extensive, it is computationally prohibitive, memory intensive, and risks catastrophic forgetting. Parameter-efficient fine-tuning (PEFT) techniques address these limitations by updating only a small subset of parameters while preserving the model’s inherent capabilities (Liu et al., 2022; Ding et al., 2023). These methods such as adapter tuning, prefix tuning, prompt tuning, and low-rank adaptation strike a balance between task-specific performance and resource efficiency, enabling cost-effective deployment of LLMs across diverse applications. We briefly introduce these them as follows.
@@ 3.1 Conceptual framework
    Excellent reasoning and planning abilities: CoT prompting and recursive self-improvement enable LLMs to decompose complex tasks into multi-step reasoning processes with verifiable intermediate states. In addition, LLMs can perform complex tasks by planning to use tools, integrate external knowledge, and learn from demonstrations.
@@
As Fig. 1 illustrates, the LLM4TR framework shapes the four-step transportation management cycle through an LLM-centric lens:
@@ 3.2 Taxonomy
Figure 2: Taxonomy and classification principles in this survey. Following this procedure, we classify the literature according to four primary roles: Information Processors, Knowledge Encoders, Component Generators, and Decision Facilitators. Binary decision rules used to assign each method to exactly one primary role: we begin by asking whether the model directly optimizes or outputs the final target (D. Decision Facilitator), which is further split into three subcategories according to the output type. If instead the focus is on whether LLMs are used as encoders or decoders. If encoders, models that consume raw, unstructured inputs and process system information occupy A. Information Processor. Next, it is classified into three detailed roles according to the content of the input information. If not the case the focus is on formalizing that information into structured, domain-specific knowledge (explicit vs. implicit formalization), it is classified under B. Knowledge Encoder. Finally, if LLMs as decoders synthesize new artifacts such as scenario simulations or network designs, it proceeds to C. Component Generator (differentiated by "before", "During" and “After” synthesis).
@@
      Example: Designing reward functions for traffic control agents in RL (Yu et al., 2024), assisting in synthesizing virtual driving environments (Zhao et al., 2024a), and providing feedback for the refinement of model component (Tian et al., 2024).
@@
Following this structured taxonomy, this survey classifies related literature using some principles shown in Fig. 2. Specifically, we classify each method according to the binary rules in Fig. 2 and the following criteria:
  1. 1.
    Input Type:
    • (a)
      Information Processors consume raw or minimally preprocessed data (e.g. sensor streams, traffic logs, and user queries) and transform it into actionable formats for downstream processing.
    • (b)
      Knowledge Encoders ingest processed representations (e.g., embeddings or metadata) to build domain-specific knowledge.
  2. 2.
    Output Form:
    • (a)
      Component Generators produce new artifacts such as textual description, synthetic data, code scripts, or model architectures that are tailored to downstream tasks.
    • (b)
      Decision Facilitators generate actionable recommendations or predictions directly supporting operational decisions.
  3. 3.
    Functional Objective:
    • (a)
      Processing roles (Processors & Encoders) emphasize data understanding and knowledge structuring.
    • (b)
      Generation roles (Generators & Facilitators) emphasize content synthesis and action guidance.
  4. 4.
    Pipeline Position:
    • (a)
      Processors appear at the front end of data pipelines.
    • (b)
      Encoders capture and formalize insights for reuse after data processing.
    • (c)
      Generators are placed within the model creation or solution design loops.
    • (d)
      Facilitators locate at the final decision execution stage.
Please note that in some studies the role of LLMs are multifaceted. For example, they may serve simultaneously as both a context encoder and a decision maker. In our classification, we have identified and assigned each method to the role that best reflects its primary functional contribution, ensuring clarity and consistency.
@@ 3.3 Research trend visualization
As an intuitive overview of the current research trend and focus, we visualize the statistics of selected papers according to our taxonomy in Fig. 3. Several key findings in these figures can be summarized as follows: (1) For the four roles defined in this paper, the "decision facilitator" attracts the most interest in existing studies, where most of them focus on managing and learning tasks such as traffic prediction, autonomous driving, and traffic signal control. This is also one of the most crowded tracks in transportation research. (2) The "component generator" is also a popular direction in which the generative capabilities of LLMs can assist in designing modules or systems. (3) Some areas are still unexplored. For example, adopt LLMs as decision makers in traffic modeling and designers for learning tasks. The former can automate the traffic simulation procedure by acting as a decision engine, while the latter can inspire the neural architecture design for spatial-temporal predictive learning. (4) From another facet, different tasks highlight different functionalities of LLMs. For example, the sensing task populates the integration of MLLMs for information fusion; Among these four tasks, the application of the "knowledge encoder" accounts for the highest proportion in learning. These intuitive visual cues can guide future research choices and systematic overviews of this research field.
@@ 4 LLMs as Information Processors
Due to the ICL capability (Brown et al., 2020) and the multimodal extension, LLMs have emerged as versatile tools for processing heterogeneous transportation data, enabling efficient encoding, analysis, and fusion of multimodal information generated by traffic participants. As can be seen in Fig. 4, this section explores three key methodological paradigms: (1) contextual encoding of traffic scenarios and task queries, (2) analytical reasoning over structured and unstructured traffic data, and (3) multimodal integration for holistic scenario understanding. These approaches collectively address challenges in handling heterogeneous information, multi-source data, and complex semantics in transportation systems, especially crucial for sensing and learning tasks.
@@ 4.1 Context encoder
A primary application of LLMs lies in encoding textual descriptions of dynamic traffic environments or task queries into machine-readable formats and task-specific patterns. This characteristic can facilitate the problem-solving process of ITS as LLMs can translate natural language instruction to domain-specific context that is compatible with other system components. Studies such as Keysan et al. (2023), Zhao et al. (2024a), and Liao et al. (2024) demonstrate how LLMs transform free-form language descriptions of driving scenarios into encodings that can be directly consumed by downstream systems. Advanced implementations extend this capability to traffic simulation frameworks. TP-GPT (Wang et al., 2024a) generates accurate SQL queries for large-scale traffic databases and natural language interpretations by parsing user requests. It also employs a multi-agent collaboration strategy and few-shot learning to handle complex analytical tasks, such as traffic pattern recognition and privacy-preserving data interpretation. TP-GPT outperforms GPT-4 and PaLM 2 on a traffic analysis benchmark called TransQuery. ChatSUMO (Li et al., 2024a) integrates LLMs with the SUMO traffic simulator to democratize traffic simulation for non-experts. The LLM (e.g., Llama3.1) processes natural language inputs and converts them into keywords, which trigger Python scripts to fetch OpenStreetMap data, generate road networks, and configure traffic conditions. By automating code generation and data interpretation, ChatSUMO reduces scenario creation time from 15 minutes to 30 seconds while achieving 96% simulation accuracy.